Starbucks has a large, global supply chain that must efficiently supply over 17,000 stores. Although the stores might appear to be very similar, they are actually very different. Depending on the location of the store, its size, and the profile of the customers served, Starbucks management configures the store offerings to take maximum advantage of the space available and customer preferences. Starbucks' actual distribution system is much more complex, but for the purpose of our exercise let's focus on a single item that is currently distributed through five distribution centers in the United States. Our item is a logo branded coffeemaker that is sold at some of the larger retail stores. The coffeemaker has been a steady seller over the years due to its reliability and rugged construction. Starbucks does not consider this a seasonal product, but there is some variability in demand. Demand for the product over the past 18 weeks is shown in the following table. (week -1 is the week before week 1 in the table, -2 is two weeks before week 1, etc.). Management would like you to experiment with some forecasting models to determine what should be used in a new system to be implemented. The new system is programmed to use one of two forecasting models: simple moving average or exponential smoothing. 3 5 54 32 45 40 48 50 50 41 45 46 7 20 28 54 48 72 64 68 62 72 6 9 46 18 8 10 11 36 24 12 55 46 13 41 52 49 42 48 -5 -4 -3 -2-1 33 55 40 50 43 1 2 45 4 WEEK Atlanta Boston 38 56 43 62 34 25 32 35 32 34 45 44 33 44 26 42 33 42 35 4 30 25 52 44 38 41 60 9t 34 Chicago 58 20 35 48 26 63 44 20 71 Dallas 35 40 64 40 38 48 43 52 42 43 41 53 45 36 40 33 LA 248 155 247 247 195 167 196 182 208 243 282 246 196 227 252 174 240 232 Total a. Consider using a simple exponential smoothing model. In your analysis, test two alpha values, 0.2 and 0.4. When using assume that the forecast for week 1 is the past three-week average (the average demand for periods -3,-2, and -1). For the model using an alpha of 0.4, assume that the forecast for week 1 is the past five-week average. (Round your answers to 2 decimal places.) an alpha value of 0.2 ES, a 0.2
ES, a 0.2 Week ATL BOS CHI Total DAL LA 1 41.00 41.67 44.67 53.00 49,33 229.67 2 40.67 35.67 44.00 44.67 38.00 203.00 3 37.33 33.67 40.67 37.33 37.00 186.00 4 38.33 35.67 32.33 34.33 41.00 181.67 5 38.33 38.33 34,33 39.33 45.00 195.33 6 41.33 42.00 39.33 42.00 46.33 211.00 7 38.00 56.67 44.33 51.00 54.33 244.33 8 34.00 49.00 61.33 60.00 52.67 257.00 34.67 40.00 54.00 64.33 48.33 241.33 10 40.67 44.00 38.33 61.00 39.00 223.00 11 46.00 40.00 49.00 51.67 38.33 225.00 12 35.33 44.00 51.67 43.33 43.33 217.67 13 38.33 39.33 58.00 40.33 46.00 222.00 ES, a 0.4 Week ATL BOS CHI DAL LA Total 1 40 42.40 47.40 45,00 43.60 218.40 2 37.80 36.40 44.60 41.80 41.60 202.20 3 40.00 38.20 47.20 43.20 41.80 210.40 4 41.00 36.40 37.00 42.20 40.80 197.40 5 36.40 36.40 38.40 37.60 40.80 189.60 6 40.20 38.40 39.00 38.80 42.80 199.20 7 39.40 41.60 44.60 46.00 50.60 222.20 8 34.40 45.60 50.80 51.20 50.20 232.20 38.00 41.00 52.00 56.80 47.20 235.00 10 40.80 45.00 47.00 59.00 47.00 238.80 11 37.20 44.40 56.60 57.00 45.40 240.60 12 36.40 40.80 49.00 52.20 40.60 219.00 13 43.40 39.20 45.00 43.00 43.00 217.80 b. Evaluate the forecasts that would have been made over the 13 weeks using the overall (at the end of the 13 weeks) mean absolute deviation, mean absolute percent error, and tracking signal as criteria. (Round your answers to 2 decimal places. Negative values should be indicated by a minus sign.) ATL BOS CHI DAL 10.36 LA Avg of DCs ES, a 0.2 MAD 11.51 9.46 20.44 9.00 MAPE 29.52 22.69 43.91 22.86 19.86 TS 0.26 1.51 -0.38 -1.61 1.67 ES, a 0.4 MAD MAPE TS
Answer #1

a) Solution:

Cell Formula Copy to
K7 =round(average(D2:D6),2) K7:P7
K8 =round(K7+(D7-K7)*0.2,2) K8:P19
R7 =ABS(D7-K7) R7:W19
Y7 =R7/D7*100 Y7:AD19
AF7 =D7-K7 AF7:AK19
R21 =AVERAGE(R7:R19) R24:W24, (Y24:AD24,(in %))
AF20 =SUM(AF7:AF19) AF20:AK20
AF23 =AF20/R21 AF23:AK23

e a AR AC AD A We ATL BOCHI DAL LA Total 58 43 62 43 248 20 3 41 155 forecast error ES, alpha 0.2 deviation percent deviation

Cell Formula Copy to
K7 =round(average(D2:D6),2) K7:P7
K8 =round(K7+(D7-K7)*0.4,2) K8:P19
R7 =ABS(D7-K7) R7:W19
Y7 =R7/D7*100 Y7:AD19
AF7 =D7-K7 AF7:AK19
R21 =AVERAGE(R7:R19) R24:W24, (Y24:AD24,(in %))
AF20 =SUM(AF7:AF19) AF20:AK20
AF23 =AF20/R21 AF23:AK23

We ATL BOCHI DAL LA 5 43 62 58 Total 47 43 248 41 155 35 20 forecast error percent deviation ES, alpha 0.2 deviation -2 55 40

b) Solution:

D F G ATL BOS CHI DAL Avg of DCs ES, α = 0.2 ΜΑD 3 10.79 9.37 18.2 10.25 7.39 11.2 4 MAPE 31.42 28.16 47.54 24.18 16.38 29.54